# Assessing the Role of Volumetric Brain Information in Multiple Sclerosis Progression

### v1

@@ 4 Conclusion
Supplementary Materials:
Assessing the Role of Volumetric Brain Information in Multiple Sclerosis Progression
Appendix A Additional LOCO-MP and feature selection results
@@
Appendix B Cox proportional hazards modeling
@@
Figures 2 and 3 show the distribution of test set C-indices across all folds for each feature group when a penalized Cox proportional hazards model (ridge penalty) is used to predict progression. Figure 2 uses the CCDP24 outcome and Figure 3 uses the 25-foot walking outcome.
@@
Figure 2: CCDP24 Cox model results. Boxplots show the distribution of penalized Cox proportional hazards model (ridge penalty) test-set C-indices across all cross-validation folds and repeats for outcome CCDP24. The $x$-axis consists of the five feature groupings and the $y$-axis shows the test set C-indices. “Top MRI” refers to the top five DBM features. The annotated C-index represents the median value for a specific feature grouping.
@@
Figure 3: S25FW Cox model results. Boxplots show the distribution of penalized Cox proportional hazards model (ridge penalty) test-set C-indices across all cross-validation folds and repeats for outcome 25FW. The $x$-axis consists of the five feature groupings and the $y$-axis shows the test set C-indices. “Top MRI” refers to the top five DBM features. The annotated C-index represents the median value for a specific feature grouping.
@@
Figures 4 and 5 show the stability of the test set C-indices across varying numbers of top LOCO-MP features for the penalized Cox proportional hazards model (ridge penalty). Figure 4 uses the CCDP24 outcome and Figure 5 uses the 25FW outcome.
@@
Figure 4: CCDP24 Cox model stability. Boxplots show the distribution of CCDP24 test-set C-indices for the conventional + top MRI and top MRI-only Cox proportional hazard (ridge) models, plotted across varying numbers of top LOCO-MP features. The $x$-axis represents the number of top LOCO-MP features selected and the $y$-axis is the C-index.
@@
Figure 5: S25FW Cox model stability. Boxplots show the distribution of S25FW test-set C-indices for the conventional + top MRI and top MRI-only Cox proportional hazards (ridge penalty) models, plotted across varying numbers of top LOCO-MP features. The $x$-axis represents the number of top LOCO-MP features selected and the $y$-axis is the C-index.

### v2

@@ 4 Discussion
Appendix A Details of LOCO-MP for survival analysis
Following the notation of Gan et al. (2023) [28], we denote the $N\times M$ data matrix as $X$ and the outcome as $Y$. Let $(X_{i},Y_{i})$ denote the feature-outcome pair for individual $i$, where $X_{i}$ is a $M$-dimensional vector and $Y_{i}=(T_{i},C_{i})$. For every minipatch $k=1\dots K$, we sample $n$ observations $I_{k}$ and $m$ features $F_{k}$ from $X$. Denote this minipatch as $X_{I_{k},F_{k}}$. A prediction model $\hat{\mu}_{k}$ is then trained on $X_{I_{k},F_{k}}$ and predictions are generated for the remaining observations $X_{{-I_{k},F_{k}}}$, which includes features $F_{k}$ and observations outside $I_{k}$. Repeating this over all $K$ minipatches, we compute the individual feature occlusion score for feature $j$ and observation $i$:
where
$\bar{\Delta}_{j}=\frac{1}{N}\sum_{i=1}^{N}\hat{\Delta}_{ij}.$ (1)
The prediction error for survival analysis models is commonly measured as the reciprocal of the overall concordance index (C-index) of risk scores between pairs of observations. However, this metric is not an aggregation of observation-level errors, preventing its direct usage in the LOCO-MP framework. To address this, we utilize discrete hazard loss functions used in deep learning survival analysis models [41, 42, 76]. Specifically, we evenly divide the survival time scale of the study into $d$ intervals: $[t_{0},t_{1}),\ldots,[t_{d-1},t_{d})$, where $t_{d}$ is the end of the study period. The patient’s event time is now denoted $T_{i}=q\;\text{if and only if}\;T_{i}\in[t_{q},t_{q+1})$.
$h(q\mid X_{i})=\mathbb{P}(T_{i}=q\mid T_{i}\geq q,X_{i}).$
$\ell_{U}(T_{i}=q)=h(q\mid X_{i})~{}S(q-1\mid X_{i}).$
$\ell_{C}(T_{i}>q)=S(q\mid X_{i}).$
Here, $\hat{\mu}_{-i}^{-j}$ encapsulates the collection of conditional hazards that are computed with the underlying minipatch model. Assuming that the patient conditional hazard probabilities are not exactly 0 or 1 over the study period, this loss framework satisfies the necessary assumptions for the theoretical guarantees of LOCO-MP [28].
LOCO-MP is advantageous over other feature selection techniques due to its ability to provide asymptotic inferential guarantees [28]. Moreover, LOCO-MP can also account for dependencies across features, a common issue in high-dimensional settings like ours – Gan et al. (2023) [28] discuss that, by generating randomly subsampled features across minipatches, LOCO-MP ensures that the predictive value of each feature is not being diminished by other strongly correlated features, since groups of strongly correlated features will not always appear in the same minipatch. The full LOCO-MP procedure and its theoretical guarantees are discussed in Gan et al. (2023) [28].
Appendix B Additional LOCO-MP and feature selection results
B.1 Full LOCO-MP results
@@
B.2 LOCO-MP under gradient boosting machines
Though we primarily study the performance of LOCO-MP under nonparametric random survival forests, the feature selection framework may utilize other machine learning survival algorithms, such as gradient boosting machines (GBM). We incorporate a survival GBM implementation [77] into LOCO-MP (GBM-LOCO), and capture its DBM feature importance ranks in OPERA I across ten 80% subsamples, analogously to the experiments in Section 3.1. We also capture the feature ranks of GBM alone, which are assigned based on the reduction in negative Cox partial log likelihood.
In Table 5, for the top six LOCO-MP random forest features, we present the ranks of their median predictive importance across ten subsamples under GBM and GBM-LOCO. The importance of the third and lateral ventricles and the cerebellum white matter is retained for both GBM and GBM-LOCO. The GBM-LOCO ranks are more concordant with RF-LOCO than are the GBM ranks, particularly for the third ventricle and precuneus. The parahippocampal region is unimportant under both GBM and GBM-LOCO. This suggests that an ensembling approach of multiple machine methods under LOCO could help further validate brain regions that are stably identified.
Region Feature Summary Median Subsample Rank
Statistic GBM GBM-LOCO
3rd Ventricle Std. Dev 6 2
Lateral Ventricle Median 1 1
Precuneus Median 18.5 7
Cerebellum White Matter Median 5 5
Lingual Median 9.5 13
Parahippocampal Median 25 21.5
Table 5: Ranks of median predictive importance across ten subsamples under GBM and GBM-LOCO. Results presented for top six LOCO-MP random survival forest features.
Appendix C Penalized Cox proportional hazards modeling
@@
Figures 13 and 14 show the distribution of test set C-indices across all folds for each feature group when a penalized Cox proportional hazards model (ridge penalty) is used to predict progression. Figure 13 uses the CCDP24 outcome and Figure 14 uses the 25-foot walking outcome. Figures 15 and 16 show the analogous plots for the Cox proportional hazards model with a lasso penalty.
@@
Figure 13: CCDP24 Cox model results (ridge). Boxplots show the distribution of penalized Cox proportional hazards model (ridge penalty) test-set C-indices across all cross-validation folds and repeats for outcome CCDP24. The $x$-axis consists of the five feature groupings and the $y$-axis shows the test set C-indices. “Top MRI” refers to the top six DBM features. The annotated C-index represents the median value for a specific feature grouping.
@@
Figure 14: S25FW Cox model results (ridge). Boxplots show the distribution of penalized Cox proportional hazards model (ridge penalty) test-set C-indices across all cross-validation folds and repeats for outcome 25FW. The $x$-axis consists of the five feature groupings and the $y$-axis shows the test set C-indices. “Top MRI” refers to the top six DBM features. The annotated C-index represents the median value for a specific feature grouping.
Figure 15: CCDP24 Cox model results (lasso). Boxplots show the distribution of penalized Cox proportional hazards model (lasso penalty) test-set C-indices across all cross-validation folds and repeats for outcome CCDP24. The $x$-axis consists of the five feature groupings and the $y$-axis shows the test set C-indices. “Top MRI” refers to the top six DBM features. The annotated C-index represents the median value for a specific feature grouping.
Figure 16: S25FW Cox model results (lasso). Boxplots show the distribution of penalized Cox proportional hazards model (lasso penalty) test-set C-indices across all cross-validation folds and repeats for outcome 25FW. The $x$-axis consists of the five feature groupings and the $y$-axis shows the test set C-indices. “Top MRI” refers to the top six DBM features. The annotated C-index represents the median value for a specific feature grouping.
@@
Figures 19 and 20 show the stability of the test set C-indices across varying numbers of top LOCO-MP features for the penalized Cox proportional hazards model (ridge penalty). Figure 19 uses the CCDP24 outcome and Figure 20 uses the 25FW outcome. Again, Figures X and Y show the analogous plots for the Cox proportional hazards model with a lasso penalty.
@@
Figure 17: CCDP24 Cox model stability (ridge). Boxplots show the distribution of CCDP24 test-set C-indices for the conventional + top MRI and top MRI-only Cox proportional hazard (ridge penalty) models, plotted across varying numbers of top LOCO-MP features. The $x$-axis represents the number of top LOCO-MP features selected and the $y$-axis is the C-index.
@@
Figure 18: S25FW Cox model stability (ridge). Boxplots show the distribution of S25FW test-set C-indices for the conventional + top MRI and top MRI-only Cox proportional hazards (ridge penalty) models, plotted across varying numbers of top LOCO-MP features. The $x$-axis represents the number of top LOCO-MP features selected and the $y$-axis is the C-index.
Figure 19: CCDP24 Cox model stability (lasso). Boxplots show the distribution of CCDP24 test-set C-indices for the conventional + top MRI and top MRI-only Cox proportional hazard (lasso penalty) models, plotted across varying numbers of top LOCO-MP features. The $x$-axis represents the number of top LOCO-MP features selected and the $y$-axis is the C-index.
Figure 20: S25FW Cox model stability (lasso). Boxplots show the distribution of S25FW test-set C-indices for the conventional + top MRI and top MRI-only Cox proportional hazards (lasso penalty) models, plotted across varying numbers of top LOCO-MP features. The $x$-axis represents the number of top LOCO-MP features selected and the $y$-axis is the C-index.
Appendix D Generalizability across study arms
To evaluate the generalizability of the LOCO-MP features, we replicated the random survival forest models in Section 3.2 using an additional internal control cohort that is demographically matched to OPERA I (OPERA II: NCT01412333) [26]. This arm was previously not used for feature selection or modeling and serves as an independent validation cohort. We use the same feature groupings and modeling process outlined in Section 2.3 to assess whether the LOCO-MP features identified with OPERA I can generalize to an unseen cohort. The RSF models are retrained on OPERA II for CCDP24 and S25FW progression with and without the conventional features, with hyperparameter tuning for number of trees. We compare models using all DBM features versus the LOCO-MP selected top DBM features, and look at median performance over thirty 80/20 data splits, analogously to the main text. Table 6 shows that the Top DBM features induce substantial C-Index gains over all DBM features for each retrained model set.
Includes Median C-Index $\Delta$
Outcome Conventional Features of Top DBM
CCDP No +0.06
CCDP Yes +0.04
S25FW No +0.15
S25FW Yes +0.12
Table 6: Median C-Index improvements when using sparse DBM feature set over all DBM features in OPERA II validation cohort ($n=340$). Results are from thirty 80/20 splits of the data.